Yekaterina Epshteyn Department of Mathematics University of Pittsburgh yee1@math.pitt.edu

Adams-Stolz Deconvolution LES Models

Abstract: We consider a family of LES models with arbitrarily high consistency error O(δ^{2N+2}) for N = 1, 2, 3.. that are based on the van Cittert deconvolution procedure. The family has been proposed and tested for LES with success by Adams and Stolz. We show that these models have an interesting and quite strong stability property. Using this property we prove an energy equality, existence, uniqueness and regularity of strong solutions and give a rigorous bound on the modeling error ∥\overline{u} − w∥ where w is the models’s solution and \overline{u} the true flow averages.